Taking the endpoints of the last interval obtained in part a) above as the initial
approximations, perform two iterations of the secant method to approximate the
root.
1
Expert's answer
2020-04-20T13:13:49-0400

As follows from question 110503, we were given an equation "f(x)=x^3-5x^2+1" and we established that its smallest positive root belongs to the interval

"[x_0,x_1]=[x_{i-1},x_i]=[0.469,0.5]."


Use the secant method to approximate the root:


"x_{i+1}=x_i-\frac{f(x_i)\cdot(x_i-x_{i-1})}{f(x_i)-f(x_{i-1})}=\\
\space\\
=0.5-\frac{-0.125\cdot(0.5-0.469)}{(-0.125)-0.0034}=0.4698."

Use this root for the second iteration, where "x_i=0.4698,x_{i-1}=0.469":


"x_{i+1}=0.4698-\frac{1.303\cdot10^{-4}\cdot(0.4698-0.469)}{1.303\cdot10^{-4}-0.0034}=0.46983."

Check the value of the function:


"f(0.46983)=9.23\cdot10^{-6}."